tfmlam:

a student in Greece discovers a pottery bowl that contains 62% of its original amount of c-14

jhonyy9:

again any details and the question ???

Narad:

The half life of C-14 is t =5570 yrs The radioactive constant is \[\lambda=\ln 2/t=\ln 2/5570= 0.0001244 \] \[m _{t}/m _{0}=0.62\] The age of the pottery is \[T=1/\lambda*\ln (1/0.62) = 1/0.0001244*\ln(1/0.62)=3842.7\] Age =3842.7years